# Reducing base drag on road vehicles using pulsed jets optimized by hybrid genetic algorithms

### v1

@@ 4.2 Base pressure analysis
Figure 7: Time-averaged streamlines for the four key control scenarios. Each row corresponds to a different scenario as detailed in Table 2: (from left to right) no actuation, steady-jet, optimal $J$, and optimal $J_{a}$. Each column represents a different vertical measurement plane at the spanwise locations of (from right to left) y/W=0,0.15,0.29,0.43,y/W=0,-0.15,-0.29,-0.43, and 0.5-0.5.
@@ 4.3 Flow field analysis
To investigate the physical mechanisms of control, a PIV analysis was performed for the four key scenarios identified in Table 2. For each case, 500 velocity fields were measured in nine vertical planes, as depicted in Figure 2. The velocity fields are normalized by the free-stream velocity, $U_{\infty}$. The mean streamlines for the five central planes of the uncontrolled baseline case are presented in Figure 7; due to the flow’s symmetry, only one side is shown. The baseline wake structure is consistent with previous studies on similar bluff bodies [46, 6, 5, 29, 28, 27], featuring a saddle point at $z/W\approx 0.5$ that separates two primary counter-rotating vortices within the main recirculation region. A smaller, wall-proximate vortex structure is also observed beneath (and downstream) the main bubble, consistent with the upward velocity induced by the lower half of the primary recirculation. The three-dimensional extent of this recirculation bubble is confined to the central portion of the base ($y/W\lesssim 0.4$), as it is no longer present in the planes aligned with the actuators ($y/W=0.43$). While the bubble’s length remains constant between the centerline and the first off-center plane ($y/W=0.15$), the core of the top vortex is displaced forward. Further outboard ($y/W=0.29$), this top vortex disappears, accompanied by a rapid decrease in the bubble’s overall length. In the outermost planes ($y/W\geq 0.43$), the flow is aligned with the freestream, though a complete analysis of this region, likely dominated by longitudinal trailing vortices [47], is precluded by the two-dimensional nature of the measurements.
@@
The effect of actuation is clearly evident in the three central PIV planes ($|y|/W=[0,0.15,0.29]$) shown in Figure 7. All three control strategies introduce a significant negative vertical velocity ($w$) into the near-wake, which profoundly alters the baseline topology. This downwash flow deflection towards the wall is highest at the steady-jet case, followed by the minimal $J_{a}$, suggesting a strong dependence on the injected mass flow. Such momentum injection significantly dampens the bottom recirculation vortex—eliminating it entirely at the centerline in the steady-jet case—while concurrently reducing the size of the top vortex. Furthermore, both the optimal $J$ and $J_{a}$ cases displace the weakened bottom vortex away from the model’s base, which is consistent with the increased base pressure suggested by the measurements in Figure 5. This localized pressure growth is maximum at the central plane, which seems to be the most affected by the actuations, following the previous flow description. This pattern of bubble compression and vertical reorganization of the recirculation is similar to that observed for the DRL-selected forcings on the identical van model [29] (in the penalized scenario). In their Cases 1–2, the DRL agent injected momentum from the lower edge so as to “anticipate” the interaction of the upper and lower shear layers, thereby moving the saddle point upstream and closer to the base and yielding a symmetric wake and high base pressure. Our optimal-$J$ actuates similarly: the center-plane PIV shows (i) suppression of the bottom vortex, (ii) contraction of the main bubble, and (iii) strengthening of the downward motion impinging on the base (Figure 7), all of which explain the pressure recovery in Figure 5. Similarly, it is highest for the steady-jet case, a condition consistent with the observed flow and pressure recovery in Figure 5. Similarly, the top vortex is brought closer to the model’s base in all three actuated cases, suggesting its involvement in the pressure drop observed previously in this region (Figure 5).
@@
An analysis of the second-order statistics, presented in Figure 9, provides further insight into the control mechanisms. The contours of streamwise velocity fluctuations ($u^{\prime}$) confirm that the most significant turbulent activity originates from the bottom jet in all actuated scenarios. Interestingly, the highest intensity of streamwise fluctuations is generated by the optimal control case (minimum $J$). In contrast, the case with the largest drag reduction (minimum $J_{a}$) trades these bottom fluctuations for a more even distribution between the bottom and bottom slots, as evidenced by the significantly higher Reynolds stresses ($u^{\prime}v^{\prime}$).
These distinct turbulent signatures are consistent with the previously discussed hypothesis. The optimal $J$ strategy appears to modify the main bubble’s shape by injecting energy with the bottom slot at a frequency that excites the natural shedding mode, generating large-scale turbulent structures. Conversely, the top jet seems to target higher-frequency, lower-energy phenomena, energizing the shear layer. Finally, the steady-jet case is distinguished by its large vertical velocity fluctuations ($v^{\prime}$), which are a direct consequence of the strong, persistent downward velocity field it imposes on the wake.
@@ 5 Conclusions
This work demonstrates an experiment-in-the-loop, model-free optimization of open-loop pulsed-jet actuation for reducing the aerodynamic base drag of a generic road-vehicle model. The optimization was carried out using a Hybrid Genetic Algorithm (HyGO), which effectively navigated a large parameter space to identify an optimal, non-intuitive control law.
The primary finding of this study is a significant and robust aerodynamic drag reduction of approximately $10\%$. This optimal control strategy was identified using a cost function carefully designed to ensure a net energy saving by simultaneously targeting drag minimization and penalizing the energy expenditure of the actuation system. The reliability of this solution was confirmed through repeated tests, which consistently reproduced the drag reduction. Crucial physical insight into the control mechanism was provided by synchronous base pressure measurements. These revealed that the actuation concentrates its effect on the lower half of the model’s base, leading to a substantial pressure recovery in this region, which is the main driver for the overall drag reduction.
A detailed analysis of the optimal control law reveals a clear differentiation in the role of the actuators. The main contributor to the drag reduction is the bottom jet, which operates at low frequencies to counteract the primary vortex shedding mechanism responsible for the bulk of the base pressure deficit. In contrast, the top jet focuses on higher-frequency, less energetic phenomena, suggesting a role in disrupting smaller-scale turbulent structures in the upper shear layer. It is also important to reflect on the architecture of the chosen cost function. The formulation of the term penalizing the injected mass flow ($J_{b}$) implies that a single duty cycle parameter for the lateral jets affects the cost function twice as much as those for the top and bottom jets. This structure may have artificially dampened the actuation from the side jets, pushing the optimizer towards solutions that minimized their contribution. Future studies could consider a modified cost function that decouples the penalty for each side actuator, which might unlock different and potentially more effective control strategies involving lateral forcing.
Finally, the most efficient solutions discovered here reproduce the same wake archetype (shorter, high-pressure base) that appears when the same van model is controlled by DRL agents trained with and without an energy term [29]. This cross-validation between a open-loop optimizer and a closed-loop, AI-based controller strengthens the generality of the proposed mechanism.
@@ Acknowledgments
The authors gratefully acknowledge Professors A. Ianiro and S. Discetti (UC3M) for their insightful comments and suggestions. The contributions of E. Amico, Professors J. Serpieri and G. Cafiero (PoliTO) in assisting with the experimental setup and van model, as well as the technical support of D. Bergmann with the wind-tunnel operation, are also sincerely appreciated.
@@
[…]
Appendix 0.A Individual evaluation process
@@
The $2.5\%$ uncertainty threshold was selected from a preliminary campaign with three distinct individuals measured ten times across the day; each measurement included both baseline and actuated segments, enabling like-for-like comparisons consistent with the optimization protocol.

### v2

@@ 4.2 Base pressure analysis
Figure 7: Time-averaged streamlines for the four key control scenarios. Each row corresponds to a different scenario as detailed in Table 2: (from left to right) no actuation, steady-jet, best-performing $J$, and best-performing $J_{a}$. Each column represents a different vertical measurement plane at the spanwise locations of (from right to left) y/W=0,0.15,0.29,0.43,y/W=0,-0.15,-0.29,-0.43, and 0.5-0.5.
@@ 4.3 Flow field analysis
To investigate the physical mechanisms of control, a PIV analysis was performed for the four key scenarios identified in Table 2. For each case, 500 velocity fields were measured in nine vertical planes, as depicted in Figure 2. The velocity fields are normalized by the free-stream velocity, $U_{\infty}$. The mean streamlines for the five central planes of the uncontrolled baseline case are presented in Figure 7; due to the flow’s symmetry, only one side is shown. The baseline wake structure is consistent with previous studies on similar bluff bodies [grandemange2013turbulent_wake_modes_bistability_flat_back, fan2020experimentalBiStableFlatBackAhmed, dalla2019simulationsBimodalFlatBackBluff, amico2024flow, amico2022deep, amico2022geneticFurgo], featuring a saddle point at $z/W\approx 0.5$ that separates two primary counter-rotating vortices within the main recirculation region. A smaller, wall-proximate vortex structure is also observed beneath (and downstream) the main bubble, consistent with the upward velocity induced by the lower half of the primary recirculation. The three-dimensional extent of this recirculation bubble is confined to the central portion of the base ($y/W\lesssim 0.4$), as it is no longer present in the planes aligned with the actuators ($y/W=0.43$). While the bubble’s length remains constant between the centerline and the first off-center plane ($y/W=0.15$), the core of the top vortex is displaced forward. Further outboard ($y/W=0.29$), this top vortex disappears, accompanied by a rapid decrease in the bubble’s overall length. In the outermost planes ($y/W\geq 0.43$), the flow is aligned with the freestream, though a complete analysis of this region, likely dominated by longitudinal trailing vortices [liu2021flowAhmed3D], is precluded by the two-dimensional nature of the measurements.
@@
The effect of actuation is clearly evident in the three central PIV planes ($|y|/W=[0,0.15,0.29]$) shown in Figure 7. All three control strategies introduce a significant negative vertical velocity ($w$) into the near-wake, which profoundly alters the baseline topology. This downwash flow deflection towards the wall is highest at the steady-jet case, followed by the minimal $J_{a}$, suggesting a strong dependence on the injected mass flow. Such momentum injection significantly dampens the bottom recirculation vortex, eliminating it entirely at the centerline in the steady-jet case, while concurrently reducing the size of the top vortex. Furthermore, both the best-performing $J$ and $J_{a}$ cases displace the weakened bottom vortex away from the model’s base, which is consistent with the increased base pressure suggested by the measurements in Figure 5. This localized pressure growth is maximum at the central plane, which seems to be the most affected by the actuations, following the previous flow description. This pattern of bubble compression and vertical reorganization of the recirculation is similar to that observed for the DRL-selected forcings on the identical van model [amico2024flow] (in the penalized scenario). In their Cases 1–2, the DRL agent injected momentum from the lower edge so as to “anticipate” the interaction of the upper and lower shear layers, thereby moving the saddle point upstream and closer to the base and yielding a symmetric wake and high base pressure. Our best-performing-$J$ actuates similarly: the center-plane PIV shows (i) suppression of the bottom vortex, (ii) contraction of the main bubble, and (iii) strengthening of the downward motion impinging on the base (Figure 7), all of which explain the pressure recovery in Figure 5. Similarly, it is highest for the steady-jet case, a condition consistent with the observed flow and pressure recovery in Figure 5. Similarly, the top vortex is brought closer to the model’s base in all three actuated cases, suggesting its involvement in the pressure drop observed previously in this region (Figure 5).
@@
An analysis of the second-order statistics, presented in Figure 9, provides further insight into the control mechanisms. The contours of streamwise velocity fluctuations ($u^{\prime}$) confirm that the most significant turbulent activity originates from the bottom jet in all actuated scenarios. Interestingly, the highest intensity of streamwise fluctuations is generated by the best-performing control case (minimum $J$). In contrast, the case with the largest drag reduction (minimum $J_{a}$) trades these bottom fluctuations for a more even distribution between the bottom and bottom slots, as evidenced by the significantly higher Reynolds stresses ($u^{\prime}v^{\prime}$).
Linking the control analysis with the associated flow-field modifications observed for the best-performing actuation, the bottom slot operates at a low frequency ($\tilde{f}_{3}\approx 0.12$), close to the natural vortex-shedding mode of the wake ($\tilde{f}_{\mathrm{sh}}\approx 0.17$ in the present configuration). Within the classical framework of bluff-body flow control, actuation in this frequency range is typically associated with interaction with the global wake instability, affecting vortex formation and base pressure [choi2008control]. Consistently, the time-averaged PIV fields reveal wake narrowing, recirculation shortening, and base-pressure recovery, which are canonical signatures of large-scale wake modification. The top slot, on the other hand, operates at a much higher frequency ($\tilde{f}_{1}/\tilde{f}_{3}\approx 7.08$), well separated from the shedding scale. According to established separation-control literature, forcing in this regime is generally associated with the response of unstable shear layers rather than with direct excitation of the global mode [GREENBLATT2000lowfreq, choi2008control]. The proximity of the top slot to the separation region and the energization of the Reynolds stresses and vorticity in the top edge of the planar fields further supports the interpretation that this actuation may influence the early shear-layer development and, indirectly, the downstream wake organization. It is emphasized, however, that in the absence of time-resolved velocity or pressure spectra, and the planar nature of the presented PIV measurements, these mechanisms cannot be isolated experimentally in the present work. The proposed interpretation should therefore be regarded as a physically consistent explanation grounded in established flow-control theory, rather than as a direct demonstration of targeted modal interactions.
@@ 5 Conclusions
This work demonstrates an experiment-in-the-loop, model-free optimization of open-loop pulsed-jet actuation for reducing the aerodynamic base drag of a generic road-vehicle model. The optimization was carried out using a Hybrid Genetic Algorithm (HyGO), which effectively navigated a large parameter space to identify a local-optimal, non-intuitive control law.
The primary finding of this study is a significant and robust aerodynamic drag reduction of approximately 8.8%. This control strategy was identified using a cost function carefully designed to ensure a net energy saving by simultaneously targeting drag minimization and penalizing the energy expenditure of the actuation system. The reliability of this solution was confirmed through repeated tests, which consistently reproduced the drag reduction. Crucial physical insight into the control mechanism was provided by synchronous base pressure measurements. These revealed that the actuation concentrates its effect on the lower half of the model’s base, leading to a substantial pressure recovery in this region, which is the main driver for the overall drag reduction.
A detailed analysis of the best-performing control law reveals a clear differentiation in the role of the actuators. The main contributor to the drag reduction is the bottom jet, which operates at low frequencies to counteract the primary vortex shedding mechanism responsible for the bulk of the base pressure deficit. In contrast, the top jet focuses on higher-frequency, less energetic phenomena, suggesting a role in disrupting smaller-scale turbulent structures in the upper shear layer. It is also important to reflect on the architecture of the chosen cost function. The formulation of the term penalizing the injected mass flow ($J_{b}$) implies that a single duty cycle parameter for the lateral jets affects the cost function twice as much as those for the top and bottom jets. This structure may have artificially dampened the actuation from the side jets, pushing the optimizer towards solutions that minimized their contribution. Future studies could consider a modified cost function that decouples the penalty for each side actuator, which might unlock different and potentially more effective control strategies involving lateral forcing.
Finally, the most efficient solutions discovered here reproduce the same wake archetype (shorter, high-pressure base) that appears when the same van model is controlled by DRL agents trained with and without an energy term [amico2024flow]. This consistency between HyGO, as an open-loop optimizer, and a DRL agent, as a closed-loop controller, AI-based controller strengthens the generality of the proposed mechanism.
@@ Acknowledgments
The authors gratefully acknowledge Professors A. Ianiro and S. Discetti (UC3M) for their insightful comments and suggestions. The contributions of E. Amico, Professors J. Serpieri and G. Cafiero (PoliTO), and Jose Luis Oterino (INTA) in assisting with the experimental setup and van model, as well as the technical support of D. Bergmann with the wind-tunnel operation, are also sincerely appreciated.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work, the authors used ChatGPT, Gemini and Grammarly in the writing process to improve the readability and language of the manuscript. After using this tool, the authors reviewed and edited the content as needed and assume full responsibility for the content of the published article.
Data availability
Data will be made available upon request.
@@
Appendix A Individual evaluation process
@@
The $2.5\%$ uncertainty threshold was established from a dedicated preliminary repeatability campaign designed to characterize the variability of the measurement process. Three representative individuals, spanning distinct regions of the actuation space, were each evaluated ten times. Each evaluation included both baseline ($D_{0}$) and actuated segments, ensuring like-for-like force comparisons consistent with the optimization protocol. Rather than relying solely on the limited number of individuals, the statistical population was constructed from the combinatorial set of pairwise discrepancies between repeated measurements. For each individual, the ten repetitions yield $45$ independent pairwise comparisons (each measurement is compared with the other 9), resulting in a total population of $135$ discrepancy values. The distribution of these discrepancies (or uncertainties) was analyzed, and the $95^{\mathrm{th}}$ percentile was found to lie at approximately $2.45\%$. This value was rounded to $2.5\%$ and adopted as the operational threshold. In statistical terms, this implies that approximately $5\%$ of evaluations are expected to exceed the threshold due to stochastic experimental variability alone, thereby triggering a third repetition rather than immediate rejection. In fact, $7.6\%$ of the evaluations required a third repetition, in close agreement with the expected confidence interval given the broader range of flow conditions explored, and notably, no individuals were ultimately discarded. This confirms that the selected threshold was sufficient to filter unreliable measurements while avoiding the elimination of genuinely high-performing solutions.